# Contemporary tensor network approaches to gapless and topological phases in an extended Bose-Hubbard ladder

### v1

@@ I.2.1 Edge mode
We show the charge and the neutral energy gaps in Fig. 6. We use the same parameters as in the main text, but $J/U=0.30$, $J/U=0.38$, and $J/U=0.46$ for the HI, the pHSF, and the SF, respectively. The left panels show the charge (upper) and the neutral (lower) energy gaps calculated under the open boundary condition in the HI (gray triangles), the pHSF (red circles), and the SF phases (blue squares). The right panels depict the results with the bias potential applied. These two different boundary conditions reveal the topological properties specific to the HI phase: The phase is gapless with the open boundary condition. However, the gapped bulk appears when the degeneracy in the edge state is broken due to the bias potential. However, the pHSF shows the same behavior as the SF; the phase is gapless regardless of the bias potential. This observation of a gapless bulk in the pHSF re-affirms the presence of superfluidity in the pHSF regime.
@@ I.2.2 Entanglement spectrum
We calculate the entanglement spectrum using the VUMPS method in Fig. 7. We plot the lowest part of the entanglement spectrum $\xi_{i},(i=1\sim 40)$ in the HI regime (Fig. 7(a)), the pHSF regime (Fig. 7(b)), and the SF regime (Fig. 7(c)). The used parameters are the same as in Fig. 6. The filled circles represent the spectrum at the infinite bond dimension $\chi\rightarrow\infty$ obtained by scaling the data with the different bond dimension $\chi=50\sim 450$. Here, we note that all the phases show the double-degeneracy in the lowest spectrum due to the two-leg ladder structure. In the HI phase, we can see the double- or four-degeneracy, and these fingerprints are well consistent with the HI in the one-leg ladder model Ejima et al. (2014), and with the Haldane phase in spin systems Pollmann et al. (2010). However, the pHSF shows similar fingerprints as the SF, and no topological properties are seen in the entanglement spectrum. This observation is in line with vanishing string order parameter, observed from the VUMPS calculation in the pHSF regime.

### v2

@@ I.2.1 Edge mode
We show the charge and the neutral energy gaps in Fig. 6. We use the same parameters as in the main text, but $J/U=0.30$, $J/U=0.38$, and $J/U=0.50$ for the HI, the pHSF, and the SF, respectively. The left panels show the charge (upper) and the neutral (lower) energy gaps calculated under the open boundary condition in the HI (gray triangles), the pHSF (red circles), and the SF phases (blue squares). The right panels depict the results with the bias potential applied. These two different boundary conditions reveal the topological properties specific to the HI phase: The phase is gapless with the open boundary condition. However, the gapped bulk appears when the degeneracy in the edge state is broken due to the bias potential. However, the pHSF shows the same behavior as the SF; the phase is gapless regardless of the bias potential. This observation of a gapless bulk in the pHSF re-affirms the presence of superfluidity in the pHSF regime.
@@ I.2.2 Entanglement spectrum
We calculate the entanglement spectrum using the VUMPS method in Fig. 7. We plot the lowest part of the entanglement spectrum $\xi_{i},(i=1\sim 40)$ in the HI regime (Fig. 7(a)), the pHSF regime (Fig. 7(b)), and the SF regime (Fig. 7(c)). The used parameters are the same as in the main text, but $J/U=0.30$, $J/U=0.38$, and $J/U=0.46$ for the HI, the pHSF, and the SF, respectively. The filled circles represent the spectrum at the infinite bond dimension $\chi\rightarrow\infty$ obtained by scaling the data with the different bond dimension $\chi=50\sim 450$. Here, we note that all the phases show the double-degeneracy in the lowest spectrum due to the two-leg ladder structure. In the HI phase, we can see the double- or four-degeneracy, and these fingerprints are well consistent with the HI in the one-leg ladder model [38], and with the Haldane phase in spin systems [21]. However, the pHSF shows similar fingerprints as the SF, and no topological properties are seen in the entanglement spectrum. This observation is in line with vanishing string order parameter, observed from the VUMPS calculation in the pHSF regime.